# Scaling Latent Reasoning via Looped Language Models

### v1

@@ 4.4 Supervised Fine-Tuning
We perform supervised fine-tuning (SFT) on a diverse corpus of approximately 8.3M examples drawn from high-quality public datasets. As shown in Table 6, our training mixture emphasizes mathematical reasoning (3.5M examples) and code generation (3.2M examples), while also incorporating scientific reasoning (808K examples) and conversational abilities (767K examples).
@@ 4.5 Reinforcement Learning Attempts
Following the SFT stage, we conducted exploratory RLVR (Reinforcement Learning with Verifiable Rewards) alignment experiments using algorithms such as DAPO [54] and GRPO [55] on the DAPO-17K dataset. However, these attempts did not yield significant performance gains over the final SFT checkpoint. We hypothesize two primary reasons for this. First, Model Saturation: the models are relatively small and had already undergone extensive SFT, which may have saturated their capabilities and left little room for further improvement via RL. Second, Infrastructure Challenges: the unique LoopLM architecture posed difficulties for our RL infrastructure. Specifically, our vLLM-based system could not efficiently perform rollouts with dynamic early exits and subsequently use that variable-depth information for the update step. This forced us to adopt a trade-off solution for training and inference.
We explored two main strategies based on these constraints:
@@
    Fixed 4-Round RL: We performed both rollouts and updates using a fixed 4 recurrent steps. In this setup, model performance increased normally but did not surpass the SFT checkpoint. Interestingly, we found that even when trained at a fixed 4-round depth, the model could still utilize fewer rounds at inference time, behaving as incentivized by the RL objective. The reason for this generalization is currently unknown.
@@
    Adaptive RL: We attempted to perform both rollouts and updates using the model’s native adaptive early exit mechanism. This approach failed to yield performance improvements, which we primarily attribute to the infrastructure challenges in handling the dynamic computational graphs.
@@ 5.4.1 Early Exit Strategies
Figure 5: Comparison of early exit strategies on MMLU. We evaluate four approaches across different average exit rounds: static baseline (red triangle), hidden state difference threshold (green squares), Ponder gate from standard pretraining (blue circles), and Ponder gate with specialized adaptive exit training from Section 3.4 (orange diamonds).
@@
We evaluate this strategy under two training configurations. The untrained configuration uses the gate as trained during our standard pretraining pipeline with the entropy-regularized objective (uniform prior KL loss). This represents the gate’s behavior when jointly optimized with language modeling throughout Stages 1-4. The trained configuration additionally applies the specialized adaptive exit loss described in Section 3.4, which explicitly teaches the gate to base stopping decisions on observed task loss improvements.
@@
    Even without specialized training, the Ponder gate from standard pretraining substantially outperforms the static baseline, validating that the entropy-regularized objective with uniform prior successfully enables adaptive computation. The gate learns to differentiate input difficulty through the general training dynamics, though it lacks the explicit supervision to correlate stopping decisions with actual performance improvements. This demonstrates that our base training approach already captures useful signals for resource allocation;
@@ 6.2 LoopLMs prevails in knowledge manipulation
Multi-hop QA tasks require huge amount of samples to learn according to [67] when training standard transformers. To study whether LoopLMs accelerate the learning process of this multi-hop knowledge manipulation task, we consider sample efficiency in learning. Specifically, we study how many different QA pairs are necessary for the trained model to achieve 100% accuracy, as well as the performance after training on a fixed budget of unique training samples. For simplicity, we focus on the task with 3-hop QA pairs. We separate all possible QA pairs into training subsets of different sizes, and compare when each model perfectly generalizes on the leave-out test set. Similarly to the Mano task, we train a standard ($6\otimes 1$) transformer as the baseline, and compare it with looped models ($6\otimes\{2,3,4\}$) to study the effect of the universal transformer. We also train an iso-flop model ($24\otimes 1$) for comparison. More details are included in Appendix B.3.
@@ 6.3 Discussion: towards understanding why LoopLM helps knowledge manipulation
Search on the parametric knowledge graph. During pretraining, language models often obtain an enormous amount of factual knowledge and learn analysis procedures with a rather shallow thinking depth. To perform more challenging tasks, the model needs to use multiple pieces of knowledge in the parameter space, which requires the model to search in-depth in the knowledge graph with directional dependencies formed by the atomic facts or knowledge. LoopLM naturally support an efficient reuse of the knowledge and algorithms stored in the parameter spaces: even though the knowledge piece is not retrieved or used in the previous calculations, the recurrent structure enables LoopLM to redo the procedure and extract necessary information.
@@ 7.2 Faithfulness
To further examine the consistency between the results of different rounds. We also compute a step-by-step agreement matrix $A$ over 1,000 Quora Question Pairs, where $A[i,j]$ counts identical labels between step $i$ and step $j$ (diagonal =1000=\!1000 by construction). See the right side of Figure 9. Adjacent steps never reach full agreement; for example, A[3,4]=361A[3,4]\!=\!361 indicates only $36.1\%$ of step-3 answers match step-4. A[2,3]=551A[2,3]\!=\!551 indicates only $44.9\%$ of step-2 answers match step-3. We also notice that when $i\geq 4$, the overlap consistency between step-$i$ and step-$i+1$, $A[i,i+1]$, is close to 1000. We think this phenomenon comes from: (1) the model does not learn to reason recursively when $i>4$. The model is trained within 4 loops; (2) as the number of loops increases, the answer gradually converges to a fixed point.
@@ Acknowledgement
We sincerely thank Zeyuan Allen-Zhu for his in-depth discussion on the physics of language model part and his enlightening insights on knowledge manipulation. We also thank Yonghui Wu, Guang Shi, Shu Zhong, Tenglong Ao, Chen Chen, Songlin Yang, Wenhao Chai, and Yuhong Chou for their insightful discussions. Special thanks to Wenjia Zhu – his words opened our eyes to what the real problems are in current models, and inspired us to explore this direction.
@@
Zixin Wen: Contributes to the theoretical analysis, Physics of LLMs experiments, paper writing, and RLVR.
@@ Appendix A Empirical Validation of Prior Choice
Figure 10: Effect of the prior over exit steps. Left: training loss (300-step sliding average) for a LoopLM with $T_{\max}=4$ under different priors on $z$. Colored curves correspond to geometric priors with parameter $\lambda\in\{0.1,\dots,0.9\}$; the red curve uses a uniform prior. Shaded regions indicate variability across runs. Right: prior probability over LoopLM steps induced by each $\lambda$ (uniform shown in red). Stronger geometric bias (larger $\lambda$) concentrates mass on shallow steps, reducing credit assignment to deeper computation.
@@
The right panel of Figure˜10 visualizes the priors. Large-$\lambda$ geometric priors concentrate mass at t=1t{=}122, starving deeper steps ($t{\geq}3$) of credit assignment; this undermines the “deeper is better” property. With a uniform prior, all depths receive comparable signal, enabling later iterations to specialize and deliver higher accuracy when maximum depth is allowed at inference.
@@
Although the uniform prior does not explicitly favor early exit, it does not preclude efficient inference: at test time we can still cap steps or apply a halting threshold. For a fixed average step budget, models trained with a uniform prior achieve a strictly better accuracy–compute Pareto frontier than those trained with geometric priors, indicating that unbiased depth exploration during pretraining turns into better deployment trade-offs.
@@ B.1 Capo: knowledge capacity
After pretraining on the bioS($N$) dataset, we assess a model’s knowledge capacity, defined as the number of bits of information it can reliably store. To make this measure comparable across models of different sizes, the raw bit count is normalized by the number of model parameters, yielding a “bits per parameter” metric. The derivation and motivation of the metric in discussed in [65]. For readers, we refer the detailed setting to Section 2.1 of [65].
@@ B.3 Multi-hop question answering on synthetic relations
We use a base model architecture with 1024 hidden dimensions, 16 attention heads, and 6 layers. We allow it to loop in $\{1,2,3,4\}$ times. Following the experimental setup in [67], we set the learning rate to 0.0005 with 1000 warmup steps and train for a total of 20,000 steps using batch size 2048. We run all experiments across 4 random seeds and report the average performance.
@@ B.5 Theory: latent thought with LoopLM
In this setting, we consider a simple single-head transformer architecture. We only use one-head and a two-layer gated MLP layer. For clearer theoretical demonstration, we use a special normalization layer $\mathrm{LN}()$ to threshold on $H$: $\mathrm{LN}(H)_{i,j}=\mathbf{1}\{H_{i,j}>0\}$. And the overall architecture for each loop is (where $Q,K,V,W_{1},W_{2}$ are all shared through layers)
$H_{i+0.5}=\mathrm{LN}(H_{i}+\mathrm{Attn}_{Q,K,V}(H_{i})),\ \mathrm{Attn}_{Q,K,V}(H_{i})=VH_{i}\text{softmax}(H_{i}^{\top}K^{\top}QH_{i})$
$H_{i+1}=\mathrm{LN}(H_{i+0.5}+W_{2}\mathrm{ReLU}(W_{1}H_{i+0.5}))$

### v2

@@ 4.4 Supervised Fine-Tuning
We perform SFT on a diverse corpus of approximately 8.3M examples drawn from high-quality public datasets. As shown in Table 6, our training mixture emphasizes mathematical reasoning (3.5M examples) and code generation (3.2M examples), while also incorporating scientific reasoning (808K examples) and conversational abilities (767K examples).
@@ 4.5 Reinforcement Learning Attempts
Following the SFT stage, we conducted exploratory RLVR (Reinforcement Learning with Verifiable Rewards) alignment experiments using DAPO [56] and GRPO [57] on the DAPO-17K dataset. These attempts did not yield significant performance gains over the final SFT checkpoint. The primary issue stemmed from the model’s dynamic early-exit mechanism. vLLM/SGLang provide fast rollouts via a fixed execution path, which breaks under LoopLM’s variable-depth computation.
We tried two approaches, neither successful:
@@
    Off-policy rollouts: We generate full four-step rollouts in vLLM, yielding four logit candidates per token. We then selected the first token to exceed the termination threshold to simulate an early exit. For updates, we used the cumulative loss up to that step, discarding later tokens and losses. This off-policy mismatch, i.e., where tokens are produced with the final depth, and losses computed at an earlier depth, did not improve performance.
@@
    Fixed 4-Round RL: To avoid off-policy issues, we performed rollouts and updates at a fixed four recurrent steps. Training progressed normally but performance did not surpass the SFT checkpoint. A like cause is scale: after having already undergone extensive SFT, these smaller models may have limited headroom for RL gains. Interestingly, the model still used fewer rounds at inference when beneficial despite being trained at four rounds. The mechanism behind this generalization remains unclear.
@@ 5.4.1 Early Exit Strategies
Figure 5: Comparison of early exit strategies on MMLU. We evaluate four approaches across different average exit rounds: static baseline (red triangle), hidden state difference threshold (green squares), Ponder gate from standard pre-training (blue circles), and Ponder gate with specialized adaptive exit training from Section 3.4 (orange diamonds).
@@
We evaluate this strategy under two training configurations. The untrained configuration uses the gate as trained during our standard pre-training pipeline with the entropy-regularized objective (uniform prior KL loss). This represents the gate’s behavior when jointly optimized with language modeling throughout Stages 1-4. The trained configuration additionally applies the specialized adaptive exit loss described in Section 3.4, which explicitly teaches the gate to base stopping decisions on observed task loss improvements.
@@
    Even without specialized training, the Ponder gate from standard pre-training substantially outperforms the static baseline, validating that the entropy-regularized objective with uniform prior successfully enables adaptive computation. The gate learns to differentiate input difficulty through the general training dynamics, though it lacks the explicit supervision to correlate stopping decisions with actual performance improvements. This demonstrates that our base training approach already captures useful signals for resource allocation;
@@ 6.2 LoopLMs prevails in knowledge manipulation
Multi-hop QA tasks require huge amount of samples to learn according to [69] when training standard transformers. To study whether LoopLMs accelerate the learning process of this multi-hop knowledge manipulation task, we consider sample efficiency in learning. Specifically, we study how many different QA pairs are necessary for the trained model to achieve 100% accuracy, as well as the performance after training on a fixed budget of unique training samples. For simplicity, we focus on the task with 3-hop QA pairs. We separate all possible QA pairs into training subsets of different sizes, and compare when each model perfectly generalizes on the leave-out test set. Similarly to the Mano task, we train a standard ($6\otimes 1$) transformer as the baseline, and compare it with looped models ($6\otimes\{2,4\}$) to study the effect of the universal transformer. We also train an iso-flop model ($24\otimes 1$) for comparison. More details are included in Appendix B.3.
@@ 6.3 Discussion: towards understanding why LoopLM helps knowledge manipulation
Search on the parametric knowledge graph. During pre-training, language models often obtain an enormous amount of factual knowledge and learn analysis procedures with a rather shallow thinking depth. To perform more challenging tasks, the model needs to use multiple pieces of knowledge in the parameter space, which requires the model to search in-depth in the knowledge graph with directional dependencies formed by the atomic facts or knowledge. LoopLM naturally support an efficient reuse of the knowledge and algorithms stored in the parameter spaces: even though the knowledge piece is not retrieved or used in the previous calculations, the recurrent structure enables LoopLM to redo the procedure and extract necessary information.
@@ 7.2 Faithfulness
To further examine the consistency between the results of different rounds. We also compute a step-by-step agreement matrix $A$ over 1,000 Quora Question Pairs, where $A[i,j]$ counts identical labels between step $i$ and step $j$ (diagonal =1000=\!1000 by construction). See the right side of Figure 9. Adjacent steps never reach full agreement; for example, A[2,4]=361A[2,4]\!=\!361 indicates only $36.1\%$ of step-2 answers match step-4. A[2,3]=551A[2,3]\!=\!551 indicates only $55.1\%$ of step-2 answers match step-3. We also notice that when $i\geq 4$, the overlap consistency between step-$i$ and step-$i+1$, $A[i,i+1]$, is close to 1000. We think this phenomenon comes from: (1) the model does not learn to reason recursively when $i>4$. The model is trained within 4 loops; (2) as the number of loops increases, the answer gradually converges to a fixed point.
@@ Acknowledgement
We sincerely thank Zeyuan Allen-Zhu for his in-depth discussion on the physics of language model part and his enlightening insights on knowledge manipulation. We thank Yuekun Yao for providing valuable insights and discussions on the multi-hop QA task. We also thank Yonghui Wu, Guang Shi, Shu Zhong, Tenglong Ao, Chen Chen, Songlin Yang, Wenhao Chai, and Yuhong Chou for their insightful discussions. Special thanks to Wenjia Zhu – his words opened our eyes to what the real problems are in current models, and inspired us to explore this direction.
@@
Zixin Wen: Contributes to the theoretical analysis, the design of adaptive exit strategies, Physics of LLMs experiments, paper writing, and RLVR.
@@ Appendix A Empirical Validation of Prior Choice
Figure 10: Effect of the prior over exit steps. Left: training loss (300-step sliding average) for a LoopLM with $T_{\max}=4$ under different priors on $z$. Colored curves correspond to geometric priors with parameter $\eta\in\{0.1,\dots,0.9\}$; the red curve uses a uniform prior. Shaded regions indicate variability across runs. Right: prior probability over LoopLM steps induced by each $\eta$ (uniform shown in red). Stronger geometric bias (larger $\eta$) concentrates mass on shallow steps, reducing credit assignment to deeper computation.
@@
The right panel of Figure˜10 visualizes the priors. Large-$\lambda$ geometric priors concentrate mass at t=1,2t{=}1,2, neglecting deeper steps ($t{\geq}3$) of credit assignment; this undermines the “deeper is better” property. With a uniform prior, all depths receive comparable signal, enabling later iterations to specialize and deliver higher accuracy when maximum depth is allowed at inference.
@@
Although the uniform prior does not explicitly favor early exit, it does not preclude efficient inference: at test time we can still cap steps or apply a halting threshold. For a fixed average step budget, models trained with a uniform prior achieve a strictly better accuracy-compute Pareto frontier than those trained with geometric priors, indicating that unbiased depth exploration during pre-training turns into better deployment trade-offs.
@@ B.1 Capo: knowledge capacity
After pre-training on the bioS($N$) dataset, we assess a model’s knowledge capacity, defined as the number of bits of information it can reliably store. To make this measure comparable across models of different sizes, the raw bit count is normalized by the number of model parameters, yielding a “bits per parameter” metric. The derivation and motivation of the metric in discussed in [67]. For readers, we refer the detailed setting to Section 2.1 of [67].
@@ B.3 Multi-hop question answering on synthetic relations
We use a base model architecture with 1024 hidden dimensions, 16 attention heads, and 6 layers. We allow it to loop in $\{1,2,4\}$ times. Following the experimental setup in [69], we set the learning rate to 0.0005 with 1000 warmup steps and train for a total of 20,000 steps using batch size 2048. We run all experiments across 4 random seeds and report the average performance.
@@ B.5 Theory: latent thought with LoopLM
In this setting, we consider a simple single-head transformer architecture. We only use one-head and a two-layer gated MLP layer. For clearer theoretical demonstration, we use a special normalization layer $\mathrm{LN}()$ to threshold on $\hat{H}$: $\mathrm{LN}(\hat{H})_{i,j}=\mathbf{1}\{\hat{H}_{i,j}>0\}$. And the overall architecture for each loop is (where $Q,K,V,W_{1},W_{2}$ are all shared through layers)
$\hat{H}_{i+0.5}=\mathrm{LN}(\hat{H}_{i}+\mathrm{Attn}_{Q,K,V}(\hat{H}_{i})),\ \mathrm{Attn}_{Q,K,V}(\hat{H}_{i})=VH_{i}\text{softmax}(\hat{H}_{i}^{\top}K^{\top}Q\hat{H}_{i})$
$\hat{H}_{i+1}=\mathrm{LN}(\hat{H}_{i+0.5}+W_{2}\mathrm{ReLU}(W_{1}\hat{H}_{i+0.5}))$